# Dual Target-Mounted RISs-Assisted ISAC Against Eavesdropping and Malicious Interference

### v1

@@ Abstract
The synergy between integrated sensing and communication (ISAC) and reconfigurable intelligent surfaces (RISs) unlocks novel applications and advanced services for next-generation wireless networks, yet also introduces new security challenges. In this study, a novel dual target-mounted RISs-assisted ISAC scheme is proposed, where a base station with ISAC capability performs sensing of two unmanned aerial vehicle (UAV) targets, one of which is legitimate and the other is eavesdropper, while communicating with the users through an RIS mounted on the legitimate UAV target. The proposed scheme addresses dual security threats posed by a hostile UAV target: eavesdropping on legitimate user communications and random interference attacks launched by a malicious RIS mounted on this eavesdropper UAV target, aiming to disrupt secure transmissions. A non-convex optimization problem maximizing the secrecy rate of the users is formulated, and a semi-definite relaxation (SDR)-based two-stage solution is developed to optimize the transmit beamforming matrix of the base station and the phase shift coefficients of the legitimate RIS. Extensive computer simulations are conducted to evaluate the robustness of the proposed solution under various system configurations. The proposed system’s communication performance is assessed using the secrecy rate metric, while the sensing performance is evaluated through the signal-to-interference-plus-noise ratio and the Cramer–Rao bound (CRB) for angle-of-departure (AoD) estimation of the eavesdropper UAV target.
Index Terms:
Integrated sensing and communication (ISAC), reconfigurable intelligent surface (RIS), secure communication, semi-definite relaxation (SDR).
@@ I Introduction
As 6G studies advance, a growing number of studies for RIS-assisted ISAC schemes have emerged in the literature [15]. In [16, 17], the sensing-capable RIS concepts have been introduced to tackle the inherent severe path attenuation of RIS-aided system designs. In [18, 19], a joint beamforming and phase optimization algorithm has been developed to satisfy quality of service (QoS) requirements of both communication users and sensing targets. Moreover, target-mounted aerial RIS-aided ISAC schemes have been presented to improve coverage [20] and facilitate accurate localization [16]. Furthermore, several studies have explored secure communication in RIS-assisted ISAC systems, considering scenarios where sensing targets could act as potential eavesdroppers intercepting legitimate users, while RISs are strategically configured to overcome security challenges effectively [21, 22]. Although prior studies have investigated the effects of eavesdropper targets in RIS-aided ISAC schemes [20, 21, 22, 23, 24], the joint impact of eavesdropper targets and malicious RIS attacks within an ISAC system has not yet been investigated.
@@
In this paper, to overcome the above challenges, a novel RIS-assisted ISAC system is proposed, where a base station (BS) equipped with a uniform linear array (ULA) performs sensing of two unmanned aerial vehicle (UAV) targets while simultaneously establishing reliable communication links with the users through an RIS mounted on the legitimate UAV. The system model considers security threats posed by an eavesdropper UAV target that intercepts wireless communication signals while a malicious RIS, mounted on the eavesdropper UAV, attempts to disrupt the users’ communication by launching random interference attacks. To investigate this dual-threat scenario, including both an eavesdropper UAV and a malicious RIS, a semi-definite relaxation (SDR)-based two-stage optimization algorithm is developed to determine the transmit beamforming matrix of the BS and the reflection coefficients of the legitimate RIS such that the system secrecy rate is maximized. The system performance is evaluated through computer simulations by analyzing the communication subsystem with the users’ secrecy rate and the sensing subsystem with the sensing signal-to-interference ratio (SINR) of UAV targets and their Cramer-Rao bounds (CRBs) for the two-dimensional (2D) angle of departures (AoDs) estimations. The major contributions of this study can be summarized as follows:
@@
  • A dual target-mounted RISs-assisted secure ISAC system is considered, where one UAV target both acts as an eavesdropper to intercept the communication between the BS and the users and launches random interference attacks using the malicious RIS to disrupt the user communication links. To remedy the eavesdropper attacks, an RIS mounted on the legitimate UAV is utilized to enhance the system secrecy rate while satisfying the sensing performance requirements.
@@
  • The optimization problem for the proposed system model is formulated to maximize the system secrecy rate while guaranteeing achievable sensing performance under the attacks of both eavesdropping and malicious RIS interference. To address this challenge, an SDR-based two-stage solution is developed to optimize both the transmit beamforming matrix of the BS and the phase shift coefficients of the legitimate RIS.
@@
  • As a communication performance metric, the sum secrecy rate of the legitimate users is investigated through extensive computer simulations. For sensing performance, the SINR of UAV targets is analyzed, and the CRB is derived to estimate the 2D AoDs of UAV targets.
@@
  • The simulation results demonstrate the effectiveness of the proposed algorithm, showing a notable performance improvement in the secrecy rate of the communication users while simultaneously enhancing key sensing performance metrics, namely SINR and CRB.
@@ II-A Dual UAV-mounted RISs-assisted ISAC System
In the proposed scheme, as illustrated in Fig. 1, a BS, equipped with $T_{x}$ transmit antennas in a ULA structure, performs sensing of two UAV targets while simultaneously establishing communication links with $K$ single-antenna legitimate users via RISs mounted on those UAVs. Among two UAV targets, one is legitimate, while the other attempts to eavesdrop on the communication users. The legitimate UAV target carries a legitimate RIS to sustain reliable communication links between the BS and the users, whereas the eavesdropping UAV target is equipped with a malicious RIS, introducing random interference attacks to disrupt the communication links between the BS and the users. Both legitimate and malicious RISs are passive with uniform planar arrays (UPAs), consisting of $N_{L}$ and $N_{M}$ reflecting elements, respectively. The reflection matrices of legitimate and malicious RIS can be respectively given as
Figure 1: Dual Target-mounted RISs-assisted ISAC scheme.
@@
$\mathbf{\Omega}_{L}\in\mathbb{C}^{N_{L}\times N_{L}}=\mathrm{diag}\left(e^{j\psi_{1}},e^{j\psi_{2}},\cdots,e^{j\psi_{N_{L}}}\right)$ (1)
@@
$\mathbf{\Omega}_{M}\in\mathbb{C}^{N_{M}\times N_{M}}=\mathrm{diag}\left(e^{j\zeta_{1}},e^{j\zeta_{2}},\cdots,e^{j\zeta_{N_{M}}}\right)$ (2)
@@
where $\psi_{l}$ is the phase shift of $l$-th reflecting element of the legitimate RIS and $\zeta_{m}$ is the phase shift of the $m$-th reflecting element of the malicious RIS, for $l\in\{1,\cdots,N_{L}\}$ and $m\in\{1,\cdots,N_{M}\}$. Here, since the malicious RIS performs random interference attacks, each $\zeta_{m}$ is assumed to be independently and identically distributed (i.i.d.) complex Gaussian random variable and follows $\sim\mathcal{CN}(0,1)$ distribution, while $\psi_{l}$ is optimized for enhancing secure transmission. In the proposed scheme, for enabling joint sensing and communication simultaneously, the BS transmits a unified signal at the $t$-th time interval as follows
$\mathbf{x}(t)\in\mathbb{C}^{T_{x}\times 1}=\mathbf{W}_{c}{\mathbf{x}_{c}}(t)+\mathbf{W}_{s}{\mathbf{x}_{s}}(t)$ (3)
$\hskip 55.48277pt=\bigl{[}\mathbf{W}_{c},\mathbf{W}_{s}\bigr{]}\bigl{[}{\mathbf{x}_{c}}(t),{\mathbf{x}_{s}}(t)\bigr{]}^{\mathrm{T}}$ (4)
@@
where $\mathbf{W}_{c}\in\mathbb{C}^{T_{x}\times K}$ and $\mathbf{W}_{s}\in\mathbb{C}^{T_{x}\times T_{x}}$ are the transmit beamforming matrix for communication users and sensing UAV targets, respectively. Here, ${\mathbf{x}_{c}}(t)\in\mathbb{C}^{K\times 1}$ is the communication signal for the legitimate users and satisfying $\mathbb{E}\{\mathbf{c}(t)\mathbf{x}_{c}(t)^{\mathrm{H}}\}=\mathbf{I}_{K}$, while and ${\mathbf{x}_{s}}(t)\in\mathbb{C}^{T_{x}\times 1}$ is frequency modulated continuous wave (FMCW) sensing signal [25] satisfying $\mathbb{E}\{\mathbf{x}_{s}(t)\mathbf{x}_{s}(t)^{\mathrm{H}}\}=\mathbf{I}_{T_{x}}$ and also orthogonal to the communication signal to prevent mutual information $\mathbb{E}\{\mathbf{x}_{s}(t)\mathbf{x}_{c}(t)^{\mathrm{H}}\}=\mathbf{0}_{T_{x}\times K}$. Therefore, for $\mathbf{W}\in\mathbb{C}^{T_{x}\times(T_{x}+K)}=\bigl{[}\mathbf{W}_{c},\mathbf{W}_{s}\bigr{]}$ being the overall transmit beamforming matrix, the covariance of the transmit signal becomes
$\mathbf{R_{x}}=\mathbb{E}\{\mathbf{x}(t)\mathbf{x}(t)^{\mathrm{H}}\}$ (5)
@@ II-B Sensing Model
In the proposed scheme, both the legitimate and eavesdropper UAV targets are assumed to have line-of-sight (LoS) links with the BS. The array responses between the BS and the UAV targets are modeled using the 2D AoD, defined by the horizontal and vertical angles $\theta^{\text{BS}}_{i}$ and $\phi^{\text{BS}}_{i}$, respectively, where $i\in\{L,E\}$ indicates legitimate UAV ($L$-UAV) and eavesdropper UAV ($E$-UAV), respectively. Then, the transmit steering vector for $i$-UAV is given by
$\mathbf{a}(\theta^{\text{BS}}_{i},\phi^{\text{BS}}_{i})\in\mathbb{C}^{1\times T_{x}}=[1,\cdots,e^{j\nu_{T}(T_{x}-1)\cos(\theta^{\text{BS}}_{i})\cos(\phi^{\text{BS}}_{i})}]$ (8)
@@
$\mathbf{Y}_{i}=\beta_{i}\mathbf{a}(\theta^{\text{BS}}_{i},\phi^{\text{BS}}_{i})^{\mathrm{H}}\mathbf{a}(\theta^{\text{BS}}_{i},\phi^{\text{BS}}_{i})\mathbf{X}+\mathbf{N}_{i}$ (9)
where $\mathbf{X}=[\mathbf{x}(1),\cdots,\mathbf{x}(L_{s})]$ and $\mathbf{N}_{i}\in\mathbb{C}^{T_{x}\times L_{s}}$ is the additive white Gaussian noise (AWGN) whose each entry follows $\mathcal{CN}(0,\sigma^{2}_{n})$ distribution. Here, $\beta_{i}$ is the complex-valued round-trip path attenuation between the BS and $i$-UAV [26], which can be given as
$\beta_{i}=\sqrt{\frac{\lambda^{2}S}{64\pi^{3}d_{i}^{4}}}$ (10)
where $S$ is the radar cross section (RCS). Therefore, the sensing SINR of the $i$-UAV target can be calculated as
$\gamma_{i}=\frac{\left\|\beta_{i}\mathbf{a}(\theta^{\text{BS}}_{i},\phi^{\text{BS}}_{i})^{\mathrm{H}}\mathbf{a}(\theta^{\text{BS}}_{i},\phi^{\text{BS}}_{i})\mathbf{W}\right\|_{\mathrm{F}}^{2}}{\left\|\beta_{j}\mathbf{a}(\theta^{\text{BS}}_{j},\phi^{\text{BS}}_{j})^{\mathrm{H}}\mathbf{a}(\theta^{\text{BS}}_{j},\phi^{\text{BS}}_{j})\mathbf{W}\right\|_{\mathrm{F}}^{2}+\sigma_{n}^{2}}$ (11)
@@
For further sensing performance evaluation, the CRB estimation of AoD pairs of legitimate and eavesdropper UAV targets is derived from the Fisher information matrix (FIM). In order to determine FIM of the $i$-UAV, first, the received echo signal given in (9) is vectorized as
$\mathbf{y}_{i}=\mathbf{p}_{i}+\mathbf{v}_{i}$ (12)
@@
$\mathbf{F}_{i}=\begin{bmatrix}\mathbf{F}_{\boldsymbol{\vartheta}_{i}\boldsymbol{\vartheta}_{i}}&\mathbf{F}_{\boldsymbol{\vartheta}_{i}\boldsymbol{\delta}_{i}}\\ \mathbf{F}_{\boldsymbol{\delta}_{i}\boldsymbol{\vartheta}_{i}}&\mathbf{F}_{\boldsymbol{\delta}_{i}\boldsymbol{\delta}_{i}}\end{bmatrix}$ (13)
@@
Here, each element of the FIM matrix in (13) can be given as
$\mathbf{F}_{i}(\epsilon,\tau)=\frac{2}{\sigma_{n}^{2}}\mathfrak{Re}\left(\frac{{d}\mathbf{p}_{i}^{\mathrm{H}}}{d\boldsymbol{\omega}_{i,\epsilon}}\frac{{d}\mathbf{p}_{i}}{d\boldsymbol{\omega}_{i,\tau}}\right)$ (15)
where $\boldsymbol{\omega}_{i,\epsilon}$ and $\boldsymbol{\omega}_{i,\tau}$ are the $\epsilon$- and $\tau$-th elements of $\boldsymbol{\omega}_{i}$ for ${\epsilon,\tau}\in\{1,2,3,4\}$. Considering the transmit steering vector of $i$-UAV target $\mathbf{a}(\theta_{i}^{\text{BS}},\phi_{i}^{\text{BS}})$ given in (8), let $\mathbf{A}_{i}=\mathbf{a}^{\mathrm{H}}(\theta_{i}^{\text{BS}}\phi_{i}^{\text{BS}})\mathbf{a}(\theta_{i}^{\text{BS}}\phi_{i}^{\text{BS}})$, while $\bar{\mathbf{A}}_{\theta_{i}^{\text{BS}}}$, $\bar{\mathbf{A}}_{\phi_{i}^{\text{BS}}}$ and $\bar{\mathbf{A}}_{\boldsymbol{\delta}_{i}}$ be the derivatives of $\mathbf{A}_{i}$ with respect to $\theta_{i}^{\text{BS}}$, $\phi_{i}^{\text{BS}}$ and $\boldsymbol{\delta}_{i}$}, respectively. Therefore, the FIM matrix elements become [27, 28]
@@
$\mathbf{F}_{\boldsymbol{\delta}_{i}\boldsymbol{\vartheta}_{i}}=\frac{2L_{s}}{\sigma_{n}^{2}}\mathfrak{Re}\left(\begin{bmatrix}{\beta_{i}^{*}\operatorname{Tr}(\mathbf{A}}_{i}\mathbf{R_{x}}(\bar{\mathbf{A}}_{\theta_{i}^{\text{BS}}})^{\mathrm{H}})\\ {\beta_{i}^{*}\operatorname{Tr}(\mathbf{A}}_{i}\mathbf{R_{x}}(\bar{\mathbf{A}}_{\phi_{i}^{\text{BS}}})^{\mathrm{H}})\end{bmatrix}[1,j]\right)$ (17)
$\mathbf{F}_{\boldsymbol{\delta}_{i}\boldsymbol{\delta}_{i}}=\frac{2L_{s}}{\sigma_{n}^{2}}\mathbf{I}_{2}\mathbf{R_{x}}{\operatorname{Tr}(\mathbf{A}}_{i}\mathbf{R_{x}}(\bar{\mathbf{A}}_{i})^{\mathrm{H}}).$ (18)

### v2

@@ Abstract
The synergy between integrated sensing and communication (ISAC) and reconfigurable intelligent surfaces (RISs) unlocks novel applications and advanced services for next-generation wireless networks, yet also introduces new security challenges. In this study, a novel dual target-mounted RISs-assisted ISAC scheme is proposed, where a base station with ISAC capability performs sensing of two unmanned aerial vehicle (UAV) targets, one of which is legitimate and the other is eavesdropper, while communicating with the users through an RIS mounted on the legitimate UAV target. The proposed scheme addresses dual security threats posed by a hostile UAV target: eavesdropping on legitimate user communications and random interference attacks launched by a malicious RIS mounted on this eavesdropper UAV target, aiming to disrupt secure transmissions. Moreover, malicious RIS interference is also optimized for a worst-case scenario, in which both the channel state information (CSI) and the transmit beamforming of the base station are assumed to be fully compromised by a malicious RIS-mounted eavesdropper UAV. A non-convex optimization problem maximizing the secrecy rate of the users is formulated, and a semi-definite relaxation (SDR)-based two-stage solution is developed to optimize the transmit beamforming matrix of the base station and the phase shift coefficients of the legitimate RIS. Extensive computer simulations are conducted to evaluate the robustness of the proposed solution under various system configurations. The proposed system’s communication performance is assessed using the secrecy rate metric, while the sensing performance is evaluated through the signal-to-interference-plus-noise ratio and the Cramer–Rao bound (CRB) for angle-of-departure (AoD) estimation of the eavesdropper UAV target.
@@ I Introduction
As 6G studies advance, a growing number of studies for RIS-assisted ISAC schemes have emerged in the literature [10]. In [26, 34], the sensing-capable RIS concepts have been introduced to tackle the inherent severe path attenuation of RIS-aided system designs. In [20, 35], a joint beamforming and phase optimization algorithm has been developed to satisfy quality of service (QoS) requirements of both communication users and sensing targets. Moreover, target-mounted aerial RIS-aided ISAC schemes have been presented to improve coverage [29] and facilitate accurate localization [6], improve air to ground communication [23], mitigate malicious jamming [5] and enhance target detection probability [4]. Furthermore, several studies have explored secure communication in RIS-assisted ISAC systems, considering scenarios where sensing targets could act as potential eavesdroppers intercepting legitimate users, while RISs are strategically configured to overcome security challenges effectively [30, 19]. Although prior studies have investigated the effects of eavesdropper targets in RIS-aided ISAC schemes [29, 30, 19, 7, 24], the joint impact of eavesdropper targets and malicious RIS attacks within an ISAC system has not yet been investigated. This critical gap motivates this study, to jointly tackle these dual threats in an ISAC system. The optimization is performed under imperfect channel state information (CSI) of the eavesdropper channel, reflecting a practically relevant scenario where the legitimate base station (BS) has incomplete and uncertain knowledge of the eavesdropper, while a malicious RIS simultaneously attempts to degrade secure transmissions through deliberate interference attacks.
I-A Motivation and Contributions
In [23, 5], the UAV-mounted RIS is employed merely as a supporting component to strengthen links for distant communication users and sensing targets, thereby improving both communication reliability and sensing accuracy within the ISAC system. Moreover, in [4], sensing vehicle targets are lodged with RISs, each equipped with amplitude and phase controllable elements to deliberately back-reflect the probing signal and enhance its detectability. In the proposed dual target-mounted RIS-assisted ISAC schemes, unlike [23, 5], which integrate RISs on external mobile vehicles to enhance the communication and sensing performance of distant users and targets, the RISs are instead mounted directly on the detectable sensing targets. However, in contrast to [4], which deploys an RIS to avoid treating the vehicle as an external target, our scheme diverges by strategically employing dual target-mounted RISs with opposing functions. While one RIS intentionally generates a malicious reflective response for legitimate users, the other is dedicated to enhancing physical layer security for those same users.
Existing RIS-assisted secure ISAC systems [29, 30, 19, 7] simply treat the sensing target as a potential eavesdropper, and an external RIS is configured solely to enhance the secrecy performance of legitimate users. In contrast, the proposed scheme addresses a far more challenging problem when one of the targets is an eavesdropper equipped with a malicious RIS that introduces unpredictable interference into the legitimate system. Although the BS and this legitimate RIS are completely unaware of the presence or behavior of the malicious RIS, the legitimate RIS is jointly optimized with the BS to reinforce the sensing and communication paths.
@@ I-A Motivation and Contributions
In this paper, to overcome the above challenges, a novel RIS-assisted ISAC system is proposed, where a BS performs sensing of two unmanned aerial vehicle (UAV) targets while simultaneously establishing reliable communication links with the users through an RIS mounted on the legitimate UAV. The system model considers security threats posed by an eavesdropper UAV target that intercepts wireless communication signals while a malicious RIS, mounted on the eavesdropper UAV, attempts to disrupt the users’ communication by launching interference attacks. The major contributions of this study can be summarized as follows:
@@
  • A dual target-mounted RISs-assisted secure ISAC system is considered, where one UAV target both acts as an eavesdropper to intercept the communication between the BS and the users and launches interference attacks using the malicious RIS to disrupt the user communication links.
  • The malicious RIS interference is assumed to encompass both random attacks and a worst-case scenario, in which the malicious RIS-mounted UAV is assumed to have perfect knowledge of the CSI and the transmit beamforming matrix of the legitimate BS.
@@
  • To mitigate the eavesdropping and malicious RIS interference threats, an optimization problem is formulated under the assumption of imperfect eavesdropper CSI at the BS. The transmit beamforming matrix of the BS and the phase shift coefficients of a legitimate RIS mounted on a legitimate UAV are optimized to maximize the system secrecy rate while ensuring that the sensing performance requirements are satisfied. To address this challenge, a semi-definite relaxation (SDR)-based two-stage solution is developed.
@@
  • As a communication performance metric, the sum secrecy rate of the legitimate users is investigated through extensive computer simulations. For sensing performance, beampattern gain, signal-to-interference-plus-noise ratio (SINR) of UAV targets is analyzed, and the Cramer-Rao bound (CRB) is derived to estimate the two-dimensional (2D) angle of departures (AoDs) of UAV targets. Moreover, a worst-case malicious RIS interference scenario is investigated, where the malicious RIS-mounted UAV eavesdropper is assumed to have perfect knowledge of the transmit beamforming matrix of the BS and the CSI of both the BS and the legitimate users.
@@
  • Through comprehensive simulation results, the effectiveness of the proposed algorithm is demonstrated, showing notable performance improvements in the secrecy rate of the communication users while simultaneously enhancing key sensing performance metrics, namely the SINR and the CRB. Furthermore, comparisons against a benchmark scheme [7] are provided to further assess the performance and robustness of the proposed approach under more challenging and demanding threat conditions.
@@ II-A Dual UAV-mounted RISs-assisted ISAC System
In the proposed scheme, as illustrated in Fig. 1, a BS, equipped with $T_{x}$ transmit antennas in a uniform linear array (ULA) structure to enable horizontal beam steering for sensing two UAV targets while simultaneously establishing communication links with $K$ single-antenna legitimate users via RISs mounted on those UAVs. Among two UAV targets, one is legitimate, while the other attempts to eavesdrop on the communication users. The legitimate UAV target carries a legitimate RIS to sustain reliable communication links between the BS and the users, whereas the eavesdropping UAV target is equipped with a malicious RIS, introducing random interference attacks to disrupt the communication links between the BS and the users. Both the legitimate and malicious RISs are assumed to be passive planar surfaces, which naturally conform to a uniform planar array (UPA) configuration, comprising $N_{L}$ and $N_{M}$ reflecting elements, respectively, to enable precise phase shift control across both horizontal and vertical dimensions.
The reflection matrices of legitimate and malicious RIS can be respectively given as
Figure 1: Target-mounted RISs-assisted ISAC scheme.
@@
$\mathbf{\Omega}_{L}\in\mathbb{C}^{N_{L}\times N_{L}}=\mathrm{diag}\left(e^{j\psi_{1}},e^{j\psi_{2}},\cdots,e^{j\psi_{N_{L}}}\right),$ (1)
@@
$\mathbf{\Omega}_{M}\in\mathbb{C}^{N_{M}\times N_{M}}=\mathrm{diag}\left(e^{j\zeta_{1}},e^{j\zeta_{2}},\cdots,e^{j\zeta_{N_{M}}}\right),$ (2)
@@
where $\psi_{l}$ is the phase shift of $l$-th reflecting element of the legitimate RIS and $\zeta_{m}$ is the phase shift of the $m$-th reflecting element of the malicious RIS, for $l\in\{1,\cdots,N_{L}\}$ and $m\in\{1,\cdots,N_{M}\}$. Since the malicious RIS performs random interference attacks, each $\zeta_{m}$ is assumed to be independently and identically distributed (i.i.d.) complex Gaussian random variable and follows $\sim\mathcal{CN}(0,1)$ distribution, while $\psi_{l}$ is optimized for enhancing secure transmission. In the proposed scheme, for enabling joint sensing and communication simultaneously, the BS transmits a unified signal at the $t$-th time interval as follows
$\mathbf{x}(t)\in\mathbb{C}^{T_{x}\times 1}=\mathbf{W}_{c}{\mathbf{x}_{c}}(t)+\mathbf{W}_{s}{\mathbf{x}_{s}}(t),$ (3)
$\hskip 55.48277pt=\bigl[\mathbf{W}_{c},\mathbf{W}_{s}\bigr]\bigl[{\mathbf{x}_{c}}(t),{\mathbf{x}_{s}}(t)\bigr]^{\mathrm{T}},$ (4)
@@
where $\mathbf{W}_{c}\in\mathbb{C}^{T_{x}\times K}$ and $\mathbf{W}_{s}\in\mathbb{C}^{T_{x}\times T_{x}}$ are the transmit beamforming matrix for communication users and sensing UAV targets, respectively. The communication signal for the legitimate users is denoted by ${\mathbf{x}_{c}}(t)\in\mathbb{C}^{K\times 1}$ and satisfying $\mathbb{E}\{\mathbf{x}_{c}(t)\mathbf{x}_{c}(t)^{\mathrm{H}}\}=\mathbf{I}_{K}$, while and ${\mathbf{x}_{s}}(t)\in\mathbb{C}^{T_{x}\times 1}$ is frequency modulated continuous wave (FMCW) sensing signal [28] satisfying $\mathbb{E}\{\mathbf{x}_{s}(t)\mathbf{x}_{s}(t)^{\mathrm{H}}\}=\mathbf{I}_{T_{x}}$ and also orthogonal to the communication signal to prevent mutual information $\mathbb{E}\{\mathbf{x}_{s}(t)\mathbf{x}_{c}(t)^{\mathrm{H}}\}=\mathbf{0}_{T_{x}\times K}$. Therefore, for $\mathbf{W}\in\mathbb{C}^{T_{x}\times(T_{x}+K)}=\bigl[\mathbf{W}_{c},\mathbf{W}_{s}\bigr]$ being the overall transmit beamforming matrix, the covariance of the transmit signal becomes
$\mathbf{R_{x}}=\mathbb{E}\{\mathbf{x}(t)\mathbf{x}(t)^{\mathrm{H}}\},$ (5)
@@ II-B Sensing Model
In the proposed scheme, both the legitimate and eavesdropper UAV targets are assumed to have line-of-sight (LoS) links with the BS. The array responses between the BS and the UAV targets are modeled using the 2D AoD, defined by the horizontal and vertical angles $\theta^{\text{BS}}_{i}$ and $\phi^{\text{BS}}_{i}$, respectively, where $i\in\{L,E\}$ indicates legitimate UAV ($L$-UAV) and eavesdropper UAV ($E$-UAV), respectively. $\theta^{\text{BS}}_{L}$ and $\phi^{\text{BS}}_{L}$ of the $L$-UAV are perfectly known to the BS; however, $\theta^{\text{BS}}_{E}$ and the angle estimates of $\phi^{\text{BS}}_{E}$ related to $E$-UAV are imperfect and mismatches may occur [13]. Therefore, these angles can be expressed as $\phi^{\text{BS}}_{E}=\tilde{\phi}^{\text{BS}}_{E}+\Delta{\phi}_{E}$ and $\theta^{\text{BS}}_{E}=\tilde{\theta}^{\text{BS}}_{E}+\Delta{\theta}_{E}$. While $\tilde{\phi}^{\text{BS}}_{E}$ and $\tilde{\theta}^{\text{BS}}_{E}$ are actual horizontal and vertical angles, and $\Delta{\phi}_{E}\sim\mathcal{CN}(0,\sigma_{\phi}^{2})$ and $\Delta{\theta}_{E}\sim\mathcal{CN}(0,\sigma_{\theta}^{2})$ are the horizontal and vertical angle errors, respectively, where $\sigma_{\theta}^{2}$ and $\sigma_{\phi}^{2}$ are the corresponding error variances.
Then, the transmit steering vector for $i$-UAV is given by
$\mathbf{a}(\theta^{\text{BS}}_{i},\phi^{\text{BS}}_{i})\in\mathbb{C}^{1\times T_{x}}=[1,\cdots,e^{j\nu_{T}(T_{x}-1)\cos(\theta^{\text{BS}}_{i})\cos(\phi^{\text{BS}}_{i})}],$ (8)
@@
$\mathbf{Y}_{i}=\beta_{i}\mathbf{a}(\theta^{\text{BS}}_{i},\phi^{\text{BS}}_{i})^{\mathrm{H}}\mathbf{a}(\theta^{\text{BS}}_{i},\phi^{\text{BS}}_{i})\mathbf{X}+\mathbf{N}_{i},$ (9)
where $\mathbf{X}=[\mathbf{x}(1),\cdots,\mathbf{x}(L_{s})]$ and $\mathbf{N}_{i}\in\mathbb{C}^{T_{x}\times L_{s}}$ is the additive white Gaussian noise (AWGN) whose each entry follows $\mathcal{CN}(0,\sigma^{2}_{n})$ distribution. $\beta_{i}$ is the complex-valued round-trip path attenuation between the BS and $i$-UAV [31], which can be given as
$\beta_{i}=\sqrt{\frac{\lambda^{2}\Lambda}{64\pi^{3}d_{i}^{4}}},$ (10)
where $\Lambda$ is the radar cross section (RCS). Therefore, the sensing SINR of the $i$-UAV target can be calculated as
$\gamma_{i}=\frac{\left\|\beta_{i}\mathbf{a}(\theta^{\text{BS}}_{i},\phi^{\text{BS}}_{i})^{\mathrm{H}}\mathbf{a}(\theta^{\text{BS}}_{i},\phi^{\text{BS}}_{i})\mathbf{W}\right\|_{\mathrm{F}}^{2}}{\left\|\beta_{j}\mathbf{a}(\theta^{\text{BS}}_{j},\phi^{\text{BS}}_{j})^{\mathrm{H}}\mathbf{a}(\theta^{\text{BS}}_{j},\phi^{\text{BS}}_{j})\mathbf{W}\right\|_{\mathrm{F}}^{2}+\sigma_{n}^{2}},$ (11)
@@
For further sensing performance evaluation, the CRB, which serves as a fundamental lower bound on the variance of any unbiased estimator [26, 13], is derived from the Fisher Information Matrix (FIM) for the AoD estimation of both the legitimate $L$-UAV and eavesdropper $E$-UAV targets. In order to determine FIM of the $i$-UAV, first, the received echo signal given in (9) is vectorized as
$\mathbf{y}_{i}=\mathbf{p}_{i}+\mathbf{v}_{i},$ (12)
@@
$\mathbf{F}_{i}=\begin{bmatrix}\mathbf{F}_{\boldsymbol{\vartheta}_{i}\boldsymbol{\vartheta}_{i}}&\mathbf{F}_{\boldsymbol{\vartheta}_{i}\boldsymbol{\delta}_{i}}\\ \mathbf{F}_{\boldsymbol{\delta}_{i}\boldsymbol{\vartheta}_{i}}&\mathbf{F}_{\boldsymbol{\delta}_{i}\boldsymbol{\delta}_{i}}\end{bmatrix}$, (13)
@@
Each element of the FIM matrix in (13) can be given as
$\mathbf{F}_{i}(\epsilon,\tau)=\frac{2}{\sigma_{n}^{2}}\mathfrak{Re}\left(\frac{{d}\mathbf{p}_{i}^{\mathrm{H}}}{d\boldsymbol{\omega}_{i,\epsilon}}\frac{{d}\mathbf{p}_{i}}{d\boldsymbol{\omega}_{i,\tau}}\right),$ (15)
where $\boldsymbol{\omega}_{i,\epsilon}$ and $\boldsymbol{\omega}_{i,\tau}$ are the $\epsilon$- and $\tau$-th elements of $\boldsymbol{\omega}_{i}$ for ${\epsilon,\tau}\in\{1,2,3,4\}$. It is worth noting that for the $E$-UAV target, since the horizontal and vertical angles are expressed as $\phi^{\text{BS}}_{E}=\tilde{\phi}^{\text{BS}}_{E}+\Delta{\phi}_{E}$ and $\theta^{\text{BS}}_{E}=\tilde{\theta}^{\text{BS}}_{E}+\Delta{\theta}_{E}$, the steering vector $\mathbf{a}(\theta_{E}^{\text{BS}}\phi_{E}^{\text{BS}})$ in (8) can be evaluated using the following trigonometric expansions:
$\cos(\tilde{\theta}^{\text{BS}}_{E}+\Delta{\theta}_{E})=\cos(\tilde{\theta}^{\text{BS}}_{E})\cos(\Delta{\theta}_{E})-\sin(\tilde{\theta}^{\text{BS}}_{E})\sin(\Delta{\theta}_{E}),$
$\cos(\tilde{\phi}^{\text{BS}}_{E}+\Delta{\phi}_{E})=\cos(\tilde{\phi}^{\text{BS}}_{E})\cos(\Delta{\phi}_{E})-\sin(\tilde{\phi}^{\text{BS}}_{E})\sin(\Delta{\phi}_{E}).$ (16)
Therefore, for $\mathbf{A}_{i}=\mathbf{a}^{\mathrm{H}}(\theta_{i}^{\text{BS}}\phi_{i}^{\text{BS}})\mathbf{a}(\theta_{i}^{\text{BS}}\phi_{i}^{\text{BS}})$, the FIM matrix elements become [15, 25]
@@
$\mathbf{F}_{\boldsymbol{\delta}_{i}\boldsymbol{\vartheta}_{i}}=\frac{2L_{s}}{\sigma_{n}^{2}}\mathfrak{Re}\left(\begin{bmatrix}{\beta_{i}^{*}\operatorname{Tr}(\mathbf{A}}_{i}\mathbf{R_{x}}(\bar{\mathbf{A}}_{\theta_{i}^{\text{BS}}})^{\mathrm{H}})\\ {\beta_{i}^{*}\operatorname{Tr}(\mathbf{A}}_{i}\mathbf{R_{x}}(\bar{\mathbf{A}}_{\phi_{i}^{\text{BS}}})^{\mathrm{H}})\end{bmatrix}[1,j]\right),$ (18)
$\mathbf{F}_{\boldsymbol{\delta}_{i}\boldsymbol{\delta}_{i}}=\frac{2L_{s}}{\sigma_{n}^{2}}\mathbf{I}_{2}\mathbf{R_{x}}{\operatorname{Tr}(\mathbf{A}}_{i}\mathbf{R_{x}}(\bar{\mathbf{A}}_{i})^{\mathrm{H}}),$ (19)
where $\bar{\mathbf{A}}_{\theta_{i}^{\text{BS}}}$, $\bar{\mathbf{A}}_{\phi_{i}^{\text{BS}}}$ and $\bar{\mathbf{A}}_{\boldsymbol{\delta}_{i}}$ be the derivatives of $\mathbf{A}_{i}$ with respect to $\theta_{i}^{\text{BS}}$, $\phi_{i}^{\text{BS}}$ and $\boldsymbol{\delta}_{i}$}, respectively.